OpenStudy (anonymous):

express y=((3x)^2)/(x-1)^2 in partial fractions. Hence, find the value of d^2y/dx^2 when x=2.

OpenStudy (anonymous):

As you can see the above link \[ \frac{9 x^2}{(x-1)^2}=\frac{9}{(x-1)^2}+\frac{18}{x-1}+9 \] Can you finish it now?